# Discriminative Sample-Guided and Parameter-Efficient Feature Space Adaptation for Cross-Domain Few-Shot Learning

### v2

@@ Abstract
[…]
@@ 1 Introduction
Existing methods typically address this challenge by first learning a set of task-agnostic (i.e., generalized) features from a large dataset. These features are then fine-tuned to a specific target task using a small training set (called support set), often comprising as few as five images per category [32, 19]. These two stages are called meta-training (or pre-training) and meta-testing, respectively. The performance of the fine-tuned model is subsequently evaluated on a separate set of test samples, the query set, where the objective is to accurately categorize each query sample into one of the classes represented in the support set. However, as illustrated in Table 1, existing methods frequently involve fine-tuning a considerable number of task-specific parameters during the meta-testing phase. This practice can lead to overfitting, especially when data is scarce [35]. Furthermore, it has been shown that shallower neural network layers generally contain more general features that can be applied to new tasks without the need for explicit fine-tuning [36, 40]. Despite this, many existing studies [19, 16, 35] attempt to adapt the entire pre-trained model to the target task, which in a limited data regime, can increase the risk of overfitting.
@@
Motivated by the above, as shown in Fig. 2, we propose to employ a set of lightweight task-specific parameters for adapting the task-agnostic feature backbone to an unseen domain. Unlike prior work [16, 35, 19, 1], we propose to employ a parameter-efficient linear transformation of features for the adaptation of a pre-trained model which significantly reduces the count of trainable parameters. Also, we propose to vary the tuning depth to fine-tune only the less transferable feature representations in the deeper layers of the network, thereby further reducing the number of trainable parameters, making it $\times3$ to 4.3×4.3$\times$4.3 × more parameter-efficient than existing methods (refer Table 1 where $d_{t}=9,7$). Moreover, we propose to utilize the sample information present within the support set to identify challenging positive examples (i.e., the most dissimilar positive examples of a given class) or challenging negative examples (i.e., the most similar negative examples for a given class) for encouraging better intra and inter-class separation during fine-tuning to encourage better cluster formation in the feature space. Consequently, as shown in Fig. 1b, our approach forms well-separated clusters in the feature space, minimizing confusing centroids and thereby enhancing query classification accuracy. We advance few-shot learning by proposing a straightforward and effective pipeline. We systematically evaluate our method on the standard cross-domain few-shot classification benchmark dataset and demonstrate superior performance to existing state-of-the-art methods.
@@ 3 Methodology
Figure 2: Illustration of our framework during meta-testing. (a): A set of task-specific parameters $(\gamma_{m},\beta_{m})\in h_{\psi}$ are attached to a ViT backbone $f_{\theta}$, up to a pre-defined tuning depth $d_{t}$, where $m=1,..,6$. (b) $h_{\psi}$ is fine-tuned on the support set using a set of learnable anchors $A_{\phi}$. (c) Query images are classified by assigning them to the nearest class centre where the fine-tuned model $f_{\hat{\theta}}$ is used to generate feature representations for the images. $L$: Number of layers in the ViT, $d_{f}$: feature fusion depth, $z_{l}$: $[cls]$ output from $l$th ViT layer, $Z$: fused feature embedding, Norm: Layer Normalization, MLP: Multi-Layer Perceptron, and MHA: Multi-head Attention.
@@ 3.2 Task-specific representation learning
Varying the tuning depth. As discussed above, recent works [19, 16, 35, 1] commonly fine-tune the full backbone for the novel task. Instead, in the limited data regime, our method varies the extent of task-specific adaption of a pre-trained model for the novel task. This approach enables customized adaptation to new tasks by selectively fine-tuning pre-trained representations across the model’s layers to meet the specific requirements of each task. Specifically, we vary the depth $d_{t}$ of task-specific parameters $h_{\psi}$ attached to the pre-trained model (i.e., the depth of layers that are fine-tuned on the target task) where $0<d_{t}<L$, and $h_{\psi}=\{\psi_{j}$, where $j=(L-d_{t}+1),...,12\}$ and $\psi_{j}=\{(\gamma_{m},\beta_{m})$, where $m=1,..,6\}$. This approach, which strategically reduces the number of tunable parameters during the meta-testing phase as detailed in Table 1, is designed to further mitigate the risk of overfitting on novel tasks.
@@ 4.1 Experimental Settings
Architecture: We limit our experiments to the standard ViT-small architecture [9]. While early ResNet-based works [19, 33] resized input images to $84\times 84$ pixels, recent ViT-based works [35, 16, 35] generally resize the images to $224\times 224$ pixels to suit the ViT architecture. Following this trend, we also employ the same image resolution. However, it is worth noting that the higher image resolution may provide the model with richer visual information, thereby potentially influencing classification performance. Therefore, we include details of backbone architectures when reporting results to facilitate a transparent comparison of methods.
@@
Fine-tuning: We experimentally determine the values for the hyperparamters, where $\delta=0.1$, $\alpha=32$, and $d_{f}=4$ give optimal performance across domains. $\gamma$ and $\beta$ parameters were initialized with constant values of one and zero, i.e., without any randomness, ensuring the reproducibility of our results. We use two separate NAdam optimizers [10] with the learning rate of $0.005$ for $h_{\psi}$ and $5$ for $A_{\phi}$. We experimentally find that this learning rate combination works well across most tasks. Fine-tuning is performed across $80$ iterations on a single Nvidia A100 GPU with 80GB. We refer to Supplementary Section LABEL:appendix:Hyperparameters for more details on hyperparameters. Finally, we report results using $d_{t}=7$ for domains seen during pre-training (in-domain) and $d_{t}=9$ for previously unseen domains (out-of-domain) as they provided the optimal average results across domains. Section 6 provides more discussion on this choice.
@@ 5.1 Comparison to state-of-the-art methods
From here onwards, we refer to our method as DIPA, DIscriminative-sample-guided and Parameter-efficient Adaptation. We evaluate DIPA with the feature extractors pre-trained under the MDL, SDL and SDL-E settings on Meta-Dataset and compare it to existing state-of-the-art methods in Table 2. To facilitate fair comparisons, given the varying meta (or pre)-training strategies and backbone architectures in prior research, we include these details alongside the results. The table is organized into two sections: one for in-domain (seen) and another for out-of-domain (unseen) dataset accuracies, including their overall averages. Additionally, we include a column for the average accuracy across commonly unseen domains, as some prior works do not report results for all the unseen domains considered in our work.
@@ 6.1 Impact of varying the fine-tuning and query classification approaches
Table 3: Comparing the impact of using the traditional NCC ($NCC_{mean}$) vs. the discriminative sample-aware loss $l_{A_{\phi}}$ for fine-tuning and $NCC_{mean}$ vs anchor-based NCC ($NCC_{A_{\phi}}$) for inference. The results are reported for the MDL setting where the first eight datasets are seen during training and the last five datasets are unseen and used for testing only. Mean accuracy and 95% confidence interval are reported.
First, we compare our fine-tuning strategy $l_{A_{\phi}}$ against the conventional mean-embedding-based NCC ($NCC_{mean}$) in Table 3 (rows A0 and A2). Here, employing $l_{A_{\phi}}$ loss for fine-tuning significantly improves classification performance, yielding up to 2.8% improvement in seen domains and 1.6% in unseen domains compared to using $NCC_{mean}$. Moreover, in terms of cluster formation, on the ImageNet dataset, $l_{A_{\phi}}$ consistently outperformed NCC in terms of inter-cluster distance, intra-cluster distance, and silhouette index by 1%, 6% and 1%, respectively.
Next, we compare our query classification strategy - $NCC_{mean}$ with another alternative. As mentioned above, during fine-tuning, DIPA learns a set of anchors $A_{\phi}$ that provide strong supervision to form well-separated and compact clusters (refer Fig. 1b). This provides an opportunity to use the fine-tuned anchors in lieu of the mean class centroids used in $NCC_{mean}$ for query classification. We denote this variant as $NCC_{A_{\phi}}$. However, while anchors $A_{\phi}$ provide strong supervisory signals for cluster formations, we observe that the anchors are placed with a small offset from the mean centroid in some cases (refer Supplementary Fig. LABEL:appendix_fig:clusters_3), suggesting that the mean class centroid which encapsulates the summary of the compact cluster might be a more representative class descriptor during inference. This hypothesis is also supported by our results, where combining $l_{A_{\phi}}$ with $NCC_{mean}$ yielded an optimal performance increase of 3.9% and 5.3% on seen and unseen domains compared to $l_{A_{\phi}}$+$NCC_{A_{\phi}}$. Therefore, we employ $A_{\phi}$ anchors for cluster formation during fine-tuning and thereafter, use the well-defined clusters to obtain mean class centroids for query classification during inference.
@@ 6.2 Impact of varying the number of tuned layers
Figure 4: Variation of accuracies for each dataset in Meta-Dataset as $d_{t}$ varies in the MDL setting. Average results are reported while more detailed results can be found in Supplementary Table LABEL:appendix_tab:tuning_depth_datasets. The dotted lines represent the relatively more challenging datasets. For each dataset, the value of $d_{t}$ that reports the highest accuracy is annotated with a dot.
Figure 5: Variation of average seen, unseen and total average accuracies across all domains in Meta-Dataset as the number of tuned layers $d_{t}$ varies in the MDL setting. Average results are reported and more detailed results can be found in Supplementary Table LABEL:appendix_tab:tuning_depth_datasets. The value of $d_{t}$ that reports the highest average for each line is annotated with a dot.
@@ 6.3 Impact of varying the feature fusion depth
We compare the performance variation for few-shot classification by varying the feature fusion depth $d_{f}$ and report the average results in Table 4. First, we note that even with $d_{f}=1$, our method outperforms existing methods by a significant margin. Next, we observe that as we increase $d_{f}$ in DIPA, the average accuracies generally show an increasing trend up to $d_{f}=4,6$, validating our choice of fusing features from multiple layers. However, the performance shows a gradual declining trend afterwards. One possible reason is that combining a higher degree of shallower layers (containing more generic patterns) will cause the model to lose focus on important domain-specific features in the deeper layers. This is also reflected in the results where $d_{f}=8,12$ report even lower accuracies than $d_{f}=1,2$.
@@
Table 4: Variation of average (Avg.) accuracies for seen, unseen and all domains as the feature fusion depth $d_{f}$ and its corresponding feature representation’s dimension ($dim$) vary in DIPA under the MDL setting. Mean results are reported and more detailed results can be found in Supplementary Table LABEL:appendix_tab:fusion-depth.
@@ 6.4 Impact of Pre-training
Table 5: The impact of varying the pre-training algorithms in SDL-E setting. Average (Avg.) results are reported and more detailed results can be found in Supplementary Table LABEL:appendix_tab:pre_training.
@@ 7 Conclusions and Limitations
In this study, we explore the efficient adaptation of deep networks for few-shot classification tasks. Our approach leverages extremely lightweight linear transformations, optimized by a discriminative sample-aware loss function, to facilitate the learning of new classes and domains with a limited number of labelled samples. This method achieves state-of-the-art performance on the challenging Meta-Dataset benchmark while ensuring parameter efficiency.

### v3

@@ Abstract
In this paper, we look at cross-domain few-shot classification which presents the challenging task of learning new classes in previously unseen domains with few labelled examples. Existing methods, though somewhat effective, encounter several limitations, which we alleviate through two significant improvements. First, we introduce a lightweight parameter-efficient adaptation strategy to address overfitting associated with fine-tuning a large number of parameters on small datasets. This strategy employs a linear transformation of pre-trained features, significantly reducing the trainable parameter count. Second, we replace the traditional nearest centroid classifier with a discriminative sample-aware loss function, enhancing the model’s sensitivity to the inter- and intra-class variances within the training set for improved clustering in feature space. Empirical evaluations on the Meta-Dataset benchmark showcase that our approach not only improves accuracy up to 7.7% and 5.3% on previously seen and unseen datasets, respectively, but also achieves the above performance while being at least 3$\sim$ 3$\times$ 3 × more parameter-efficient than existing methods, establishing a new state-of-the-art in cross-domain few-shot learning. Our code is available at https://github.com/rashindrie/DIPA.
@@ 1 Introduction
Existing methods typically address this challenge by first learning a set of task-agnostic (i.e., generalized) features from a large dataset. These features are then fine-tuned to a specific target task using a small training set (called support set), often comprising as few as five images per category [55, 37]. These two stages are called meta-training (or pre-training) and meta-testing, respectively. The performance of the fine-tuned model is subsequently evaluated on a separate set of test samples, the query set, where the objective is to accurately categorize each query sample into one of the classes represented in the support set. However, as illustrated in Table 1, existing methods frequently involve fine-tuning a considerable number of task-specific parameters during the meta-testing phase. This practice can lead to overfitting, especially when data is scarce [59]. Moreover, it has been shown that shallower neural network layers generally contain more general features that can be applied to new tasks without the need for explicit fine-tuning [60, 64]. Despite this, many existing studies [37, 28, 59] attempt to adapt the entire pre-trained model to the target task, which in a limited data regime, can increase the risk of overfitting.
@@
Motivated by the above, as shown in Fig. 2, we propose to employ a set of lightweight task-specific parameters for adapting the task-agnostic feature backbone to an unseen domain. Unlike prior work [28, 59, 37, 2], we propose to employ a parameter-efficient linear transformation of features for the adaptation of a pre-trained model which significantly reduces the count of trainable parameters. Also, we propose to vary the tuning depth to fine-tune only the less transferable feature representations in the deeper layers of the network, thereby further reducing the number of trainable parameters, making it $\times3$ to 4.3×4.3$\times$4.3 × more parameter-efficient than existing methods (refer Table 1 where $d_{t}=9,7$). Moreover, we propose to utilize the sample information present within the support set to identify challenging positive examples (i.e., the most dissimilar positive examples of a given class) or challenging negative examples (i.e., the most similar negative examples for a given class) for encouraging better intra and inter-class separation during fine-tuning to encourage better cluster formation in the feature space. Consequently, as shown in Fig. 1b, our approach forms well-separated clusters in the feature space, minimizing confusing centroids and thereby enhancing query classification accuracy. We advance few-shot learning with a straightforward and effective pipeline. We systematically evaluate our method on the standard cross-domain few-shot classification benchmark and demonstrate superior performance to all well-known methods.
@@ 3 Methodology
Figure 2: Illustration of our framework during meta-testing. (a): A set of task-specific parameters $(\gamma_{m},\beta_{m})\in h_{\psi}$ are attached to a ViT backbone $f_{\theta}$, up to a pre-defined tuning depth $d_{t}$, where $m=1,..,6$. (b) $h_{\psi}$ is fine-tuned on the support set using a set of learnable anchors $A_{\phi}$. (c) Query images are classified by assigning them to the nearest class centre using the fine-tuned model $f_{\hat{\theta}}$. $L$: Number of layers in the ViT, $d_{f}$: feature fusion depth, $z_{l}$: $[cls]$ output from $l$th ViT layer, $Z$: fused feature embedding, Norm: Layer Normalization, MLP: Multi-Layer Perceptron, and MHA: Multi-head Attention.
@@ 3.2 Task-specific representation learning
Varying the tuning depth. As discussed above, recent works [37, 28, 59, 2] commonly fine-tune the full backbone for the novel task. Instead, our method varies the extent of task-specific adaption of a pre-trained model for the novel task. This approach enables customized adaptation to new tasks by selectively fine-tuning pre-trained representations across the model’s layers to meet the specific requirements of each task. Specifically, we vary the depth $d_{t}$ of task-specific parameters $h_{\psi}$ attached to the pre-trained model (i.e., the depth of layers that are fine-tuned on the target task) where $0<d_{t}<L$, and $h_{\psi}=\{\psi_{j}$, where $j=(L-d_{t}+1),...,12\}$ and $\psi_{j}=\{(\gamma_{m},\beta_{m})$, where $m=1,..,6\}$. This approach, which strategically reduces the number of tunable parameters during the meta-testing phase as detailed in Table 1, is designed to further mitigate the risk of overfitting on novel tasks.
@@ 4.1 Experimental Settings
Architecture: We limit our experiments to the standard ViT-small architecture [18]. Recent ViT-based works [28, 59], including ours, typically resize input images to $224\times 224$ pixels, deviating from earlier approaches that use $84\times 84$ pixels [37, 56]. We acknowledge that higher resolution may impact classification performance due to richer visual information. Therefore, we include backbone architecture details in our results for transparent method comparison.
@@
Fine-tuning: We experimentally determine the values for the hyperparamters, where $\delta=0.1$, $\alpha=32$, and $d_{f}=4$ give optimal performance across domains. $\gamma$ and $\beta$ parameters were initialized with constant values of one and zero, i.e., without any randomness, ensuring the reproducibility of our results. We use two separate NAdam optimizers [19] with the learning rate of $0.005$ for $h_{\psi}$ and $5$ for $A_{\phi}$. We experimentally find that this learning rate combination works well across most tasks. Fine-tuning is performed across $80$ iterations on a single Nvidia A100 GPU with 80GB. We refer to Supplementary Section 10.3 for more details on hyperparameters. Finally, we report results using $d_{t}=7$ for domains seen during pre-training (in-domain) and $d_{t}=9$ for previously unseen domains (out-of-domain) as they provided the optimal average results across domains. Section 6 provides more discussion on this choice.
@@ 5.1 Comparison to state-of-the-art methods
From here onwards, we refer to our method as DIPA, DIscriminative-sample-guided and Parameter-efficient feature Adaptation. We evaluate DIPA with the feature extractors pre-trained under the MDL, SDL and SDL-E settings on Meta-Dataset and compare it to existing state-of-the-art methods in Table 2. To facilitate fair comparisons, given the varying meta (or pre)-training strategies and backbone architectures in prior research, we include these details alongside the results. The table is organized into two sections: one for in-domain (seen) and another for out-of-domain (unseen) dataset accuracies, including their overall averages. Additionally, we include a column for the average accuracy across commonly unseen domains, as some prior works do not report results for all the unseen domains considered in our work.
@@ 6.1 Impact of varying the fine-tuning and query classification approaches
Table 3: Comparing the impact of using the traditional NCC ($NCC_{mean}$) vs. $l_{A_{\phi}}$ loss for fine-tuning and $NCC_{mean}$ vs anchor-based NCC ($NCC_{A_{\phi}}$) for inference. The results are reported for the MDL setting where the first eight datasets are seen during training and the last five datasets are unseen and used for testing only. Mean accuracy and 95% confidence interval are reported.
First, we compare our fine-tuning strategy $l_{A_{\phi}}$ against the conventional mean-embedding-based NCC ($NCC_{mean}$) in Table 3 (rows A0 and A2). Here, employing $l_{A_{\phi}}$ for fine-tuning significantly improves classification performance, yielding up to 2.8% improvement in seen domains and 1.6% in unseen domains compared to using $NCC_{mean}$. Moreover, in terms of cluster formation, on the ImageNet dataset, $l_{A_{\phi}}$ consistently outperformed $NCC_{mean}$ in terms of inter-cluster distance, intra-cluster distance, and silhouette index by 1%, 6% and 1%, respectively.
Next, we compare our query classification strategy - $NCC_{mean}$ with another alternative. As mentioned above, during fine-tuning, DIPA learns a set of anchors $A_{\phi}$ that provide strong supervision to form well-separated and compact clusters (refer Fig. 1b). This provides an opportunity to use the fine-tuned anchors in lieu of the mean class centroids used in $NCC_{mean}$ for query classification. We denote this variant as $NCC_{A_{\phi}}$. However, while anchors $A_{\phi}$ provide strong supervisory signals for cluster formations, we observe that the anchors are placed with a small offset from the mean centroid in some cases (refer Supplementary Fig. 8), suggesting that the mean class centroid which encapsulates the summary of the compact cluster might be a more representative class descriptor during inference. This hypothesis is also supported by our results, where combining $l_{A_{\phi}}$ with $NCC_{mean}$ yielded an optimal performance increase of 3.9% and 5.3% on seen and unseen domains compared to $l_{A_{\phi}}$+$NCC_{A_{\phi}}$. Therefore, we employ $A_{\phi}$ anchors for cluster formation during fine-tuning and thereafter, use the well-defined clusters to obtain mean class centroids for query classification during inference.
@@ 6.2 Impact of varying the number of tuned layers
Figure 4: Variation of accuracies for each dataset in Meta-Dataset as $d_{t}$ varies in the MDL setting. Average results are reported while more detailed results can be found in Supplementary Table 8. The dotted lines represent the relatively more challenging datasets. For each dataset, the value of $d_{t}$ that reports the highest accuracy is annotated with a dot.
Figure 5: Variation of average seen, unseen and total average accuracies across all domains in Meta-Dataset as the number of tuned layers $d_{t}$ varies in the MDL setting. Average results are reported and more detailed results can be found in Supplementary Table 8. The value of $d_{t}$ that reports the highest average for each line is annotated with a dot.
@@ 6.3 Impact of varying the feature fusion depth
We compare the performance variation for few-shot classification by varying the feature fusion depth $d_{f}$ and report the average results in Table 4. First, we note that even without feature fusion (when $d_{f}=1$), our method outperforms existing methods by a significant margin. Next, we observe that as we increase $d_{f}$ in DIPA, the average accuracies generally show an increasing trend up to $d_{f}=4,6$, validating our choice of fusing features from multiple layers. However, the performance shows a gradual declining trend afterwards. One possible reason is that combining a higher degree of shallower layers (containing more generic patterns) will cause the model to lose focus on important domain-specific features in the deeper layers. This is also reflected in the results where $d_{f}=8,12$ report even lower accuracies than $d_{f}=1,2$.
@@
Table 4: Variation of average (Avg.) accuracies for seen, unseen and all domains as the feature fusion depth $d_{f}$ and its corresponding feature representation’s dimension ($dim$) under the MDL setting. Mean results are reported and more detailed results can be found in Supplementary Table 9.
@@ 6.4 Impact of Pre-training
Table 5: The impact of varying the pre-training algorithms in SDL-E setting. Average (Avg.) results are reported and more detailed results can be found in Supplementary Table 10.
@@ 7 Conclusions and Limitations
In this study, we explore the efficient adaptation of neural networks for few-shot classification. Our approach leverages extremely lightweight linear transformations, optimized by a discriminative sample-aware loss function, to learn new classes and domains with a limited number of labelled samples. This method achieves state-of-the-art performance on the challenging Meta-Dataset benchmark while ensuring parameter efficiency.